OpenStudy (simmy2497):

pleaseeeeeee help me !!!!!! Write the point-slope form of the equation of the line that passes through the point (1, 3) and has a slope of 2. Include your work in your final answer. Type your answer in the box provided or use the upload option to submit your solution.

jimthompson5910 (jim_thompson5910):

Does this formula look familiar \[\Large y-y_1 = m(x-x_1)\]

OpenStudy (simmy2497):

not really

jimthompson5910 (jim_thompson5910):

so your teacher hasn't gone over point slope form at all?

OpenStudy (simmy2497):

she has but its just really confusing me

jimthompson5910 (jim_thompson5910):

you're more familiar with y = mx+b maybe?

OpenStudy (simmy2497):

kinda

jimthompson5910 (jim_thompson5910):

those two forms are different ways to express linear equations where m = slope with point slope form (x1,y1) is the point the line goes through

jimthompson5910 (jim_thompson5910):

So all we do is plug in the given info m = 2 x1 = 1 y1 = 3 to get... \[\Large y - y_1 = m(x-x_1)\] \[\Large y - 3 = 2(x-1)\] and that's all there is to it

OpenStudy (simmy2497):

oh ok that kinda makes more sense.!! y−3=2(x−1) would be the answer?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

you just need to memorize that basic template. Then make replacements

OpenStudy (simmy2497):

Thank you!

jimthompson5910 (jim_thompson5910):

you're welcome